# Closed Orbits of Dynamically Convex Reeb Flows: Towards the HZ- and Multiplicity Conjectures

### v2

@@ 4.1.1. Proof of Part (M1) of Theorem A
Next, let $x=x_{0},\ldots,x_{r}$ be all prime closed orbits which are eventually ${\mathbb{Q}}$-visible, i.e., $x_{j}^{k}\in{\mathcal{P}}$ for some $k\in{\mathbb{N}}$. Note that the orbits $x_{j\geq 1}$ might be ${\mathbb{Q}}$-invisible and/or have action below ${\mathcal{A}}(x)$. In fact, $x_{j}^{s}$ with $j\geq 1$ is necessarily ${\mathbb{Q}}$-invisible when ${\mathcal{A}}\big{(}x_{j}^{s}\big{)}\leq{\mathcal{A}}(x)$. For each $j\geq 1$, we denote by $s_{j}$ the largest such $s$, setting $s_{j}=0$ if ${\mathcal{A}}(x_{j})>{\mathcal{A}}(x)$. Our goal is to show that there are at least $n$ closed orbits $x_{j}$, i.e., $r+1\geq n$.
@@
Indeed, then the only closed orbits with degree in ${\mathcal{I}}$ must be from $\Gamma_{0}$. There are exactly $n$ integer spots in ${\mathcal{I}}$ of parity $n+1$, which must all be filled.
@@
Here in the first equality we used Lemma 2.8. The second inequality follows from the fact that $\operatorname{CH}(x)$ is concentrated in degree $n+1$ by Corollary D. As a consequence, $x^{k}\in{\mathcal{P}}$. We will need the following observation.
@@
We already know that $\operatorname{CH}(x;{\mathbb{Q}})$ and $\operatorname{CH}\big{(}x^{k};{\mathbb{Q}}\big{)}$ are supported in one degree and one-dimensional. Hence we only need to determine the degree shift $L$. Let $\varphi$ be the Poincaré return map of $x$. Then the return map of $x^{l}$ is $\varphi^{l}$. By (2.10) and (2.16),
$1=\dim\operatorname{CH}\big{(}x^{l};{\mathbb{Q}}\big{)}\leq\dim\operatorname{HF}\big{(}\varphi^{l};{\mathbb{Q}}\big{)}$
for all $l\in{\mathbb{N}}$. Moreover, $\operatorname{CH}(x;{\mathbb{Q}})=\operatorname{HF}(\varphi;{\mathbb{Q}})$ due to [Fe20, Thm. 1.2]; see also [Fe21]. Furthermore, as readily follows from [GG10, Thm. 1.1] (and its proof),
@@
since $k$ is admissible, with the shift $L$ exactly as in (D1) and (D2) depending on whether $x$ is totally degenerate or not. Combining these facts we conclude that $\operatorname{HF}\big{(}\varphi^{k};{\mathbb{Q}}\big{)}$ is one-dimensional and supported in $\deg(x^{k})$, and thus $\deg(x^{k})=\deg(x)+L$. ∎
@@
$\begin{split}\deg\big{(}x^{k_{0}-1}\big{)}&=(k_{0}-1)\operatorname{\hat{\mu}}(x)+q\\ &\leq\operatorname{\hat{\mu}}\big{(}x^{k_{0}}\big{)}-\operatorname{\hat{\mu}}(x)+q\\ &\leq\operatorname{\hat{\mu}}\big{(}x^{k_{0}}\big{)}-(n+1)\\ &\leq d+\eta-(n+1)\\ &<d-n.\end{split}$
(In fact, in this case $d=\operatorname{\hat{\mu}}\big{(}x^{k_{0}}\big{)}$ because $x$ is totally degenerate.) Thus $\deg\big{(}x^{k}\big{)}<d-n$ proving (4.4) when $x$ is totally degenerate.
@@
This guarantees, again by (4.2), that $x_{j}^{k_{j}-s}$ is an admissible odd iterate of $x_{j}^{s}$. Finally, by Lemma 2.8
@@
This completes the proof of (4.3) and the proof of Part (M1).
4.1.2. Proof of Part (M2)
This part is now almost obvious and, in fact, not completely new; cf. Corollary 3.7 and [GG20, GK, LZ]. Due to non-degeneracy, a closed orbit is ${\mathbb{Q}}$-visible if and only if it is good. The parity part follows immediately from Part (O1) of Theorem E since now the degree is equal to the Conley–Zehnder index. Then, keeping the notation from the proof of Part (M1), the fact that there are exactly $n$ non-alternating orbits follows from that the interval ${\mathcal{I}}$ has exactly $n$ spots of parity $n+1$ and that distinct good orbits have distinct indices by again (O1) and that $N$ is even.
@@ 4.1.3. Proof of Theorem B
Here is a detailed argument. Recall that when $M$ is centrally symmetric, there are two types of closed Reeb orbits in $M$: symmetric and asymmetric. The former are the closed orbits $x$ such that $-x(t)=x(t+T/2)$ where $T={\mathcal{A}}(x)$ is the period of $x$, and the latter are the orbits $x$ such that $-x$ and $x$ are geometrically distinct. Clearly, asymmetric orbits come in pairs, $(x,-x)$, with both $x$ and $-x$ having the same action and the same index. Thus, by Corollary D, the flow has no asymmetric orbits.
To finish the proof it suffices to show that all symmetric closed orbits are non-alternating. This is a standard observation going back in some form to [LLZ] or even earlier; below we closely follow [GiM].
Consider the homogeneous of degree two Hamiltonian $H$ on ${\mathbb{R}}^{2n}$ with $\{H=1\}=M$. The flow of $H$ on $M$ agrees with the Reeb flow on $M$. Let $\Psi^{t}$ be the linearized flow of $H$ in ${\mathbb{R}}^{2n}$ along a closed symmetric orbit $x$ in $M$ with period $T$. Using the constant trivialization $T{\mathbb{R}}^{2n}\cong{\mathbb{R}}^{2n}\times{\mathbb{R}}^{2n}$ we can view $\Psi^{t}$ as a path in $\operatorname{Sp}(2n)$. Then $\Psi^{T}=(\Psi^{T/2})^{2}$, since the trivialization and $H$ are both centrally symmetric, and therefore, $\Psi^{T}$ is non-alternating. To be more precise, $\Psi^{T}$ has an even number of eigenvalues in the range $(-1,0)$ and $-1$ is not an eigenvalue since the Reeb flow is non-degenerate. Fix a point $z$ on $x$ and consider the decomposition $T_{z}{\mathbb{R}}^{2n}=\xi_{z}\oplus{\mathbb{R}}^{2}$, where $\xi_{z}\subset T_{z}M$ is the contact plane and the second factor is the span of the Reeb vector field vector at $z$ and the radial direction. The decomposition is invariant under $\Psi^{T}\colon T_{z}{\mathbb{R}}^{2n}\to T_{z}{\mathbb{R}}^{2n}$, which then decomposes as $\Phi\oplus{\mathit{i}d}$, where $\Phi\colon\xi_{z}\to\xi_{z}$ is the linearized Reeb flow in time $T$. It follows that $\Phi$ and $\Psi^{T}$ have the same eigenvalues other than 1. Therefore, $\Phi$ and hence, by definition, $x$ are also non-alternating.
@@
It is illuminating to examine the role of symmetry, non-degeneracy and our other main results in the proof of Theorem B. The non-degeneracy condition is used to have non-alternating closed orbits defined and then, in some sense even more crucially, to guarantee that all iterates except the bad ones are ${\mathbb{Q}}$-visible; cf. Examples 2.12 and 2.13. The lower bound $n$ on the number of non-alternating prime orbits follows in essence from [GK, LZ]; cf. Corollary 3.7. The genuinely new part of Theorem A is not needed here. However, we would have the same lower bound on the total number of prime closed orbits even without non-degeneracy by (M1). Ditto, if we replaced dynamical convexity by convexity or strong dynamical convexity in Theorem B, which are admittedly more restrictive conditions, by [LLZ] or [GiM]; cf. Corollary 3.8.
The symmetry condition together with the fact that distinct orbits have distinct actions (Corollary D) is central to the proof of Theorem B to rule out the existence of asymmetric closed orbits and this step is independent of the non-degeneracy requirement.
@@ 4.2. Proof of Theorem E
The proof relies on Theorem C and Corollary D, which are the central parts of the argument. We start with the proof of Part (O1), noting first that the existence of $x\in{\mathcal{P}}$ with $\deg(x)=m$ for any $m\in n-1+2{\mathbb{N}}$ is an immediate consequence of (2.7) and (2.19).
@@
It immediately follows from (4.9) and (4.10) that sequence $\deg(I_{i})$ has the same parity. Since $\deg(I_{0})=n$, all $\deg(I_{i})$ have parity $n$.
@@
Furthermore, if we had $y\in{\mathcal{P}}$ with $q:=\deg(y)\in n+2{\mathbb{N}}$ we would have at least two orbits in ${\mathcal{P}}$ of degree $q-1$ or $q+1$ due to (2.19) or, more specifically, (2.20) and Theorem C. This is impossible since, as we have shown, for every $m\in n-1+2{\mathbb{N}}$ an orbit $x\in{\mathcal{P}}$ with $\deg(x)=m$ is unique. This completes the proof of the statement that the degree map $\deg$ on ${\mathcal{P}}$ takes values in $n-1+2{\mathbb{N}}$ and is a bijection.
@@
The statement that the ordering of the orbits in ${\mathcal{P}}$ (or the bars in ${\mathcal{B}}$) by the action agrees with the ordering by the index is simply a reformulation of the fact that the sequence $\deg(x_{i})$ is strictly monotone increasing. (Recall that the sequence $x_{i}$ is originally ordered by the action.) Alternatively, this statement is also a consequence of [GG20, Thm. 1.3]. This concludes the proof of Part (O1).
@@
When the flow is non-degenerate, Part (O1) has an analogue for any field ${\mathbb{F}}$ with $\deg(x)$ replaced my $\mu(x)$, although the statement is then less precise. Namely, the sequences are $\mu(x_{i})$ and $\deg(I_{i})$ are still increasing although now not necessarily strictly. However, for every $m\in n-1+2{\mathbb{N}}$ we still have exactly one ${\mathbb{F}}$-visible closed orbit $x$ with $\mu(x)=m$ and all $\deg(I_{i})$ have parity $n$. The key difference is that in this case we cannot rule out the existence of closed ${\mathbb{F}}$-visible orbits $x$ with $\mu(x)$ of parity $n$ and intervals in the sequences $\deg(x_{i})$ and $\deg(I_{i})$ where these sequences are constant. The proof is a subset of the proof of Part (O1) combined with the fact that $\operatorname{supp}\operatorname{SH}(x;{\mathbb{F}})=\{\mu(x),\mu(x)+1\}$ for any field ${\mathbb{F}}$ when $x$ is non-degenerate, i.e., the second part of Corollary D holds automatically. We do not know to what degree Part (O2) extends to this case.
@@ 4.3.1. Generalities
This proposition is proved in Section 4.3.2. Here we only point out that this is the only part of the proof where the finiteness and dynamical convexity conditions are used. The rest of the argument is independent of these requirements. Note also that the sequence $C_{l}$ can be taken independent of the field ${\mathbb{F}}$. However, we do not need this fact.
@@
Thus, fix $t\in{\mathbb{R}}_{+}\setminus{\mathcal{S}}(\alpha)$ and let $t^{\prime}>t$ be such that $[t,t^{\prime}]\cap{\mathcal{S}}(\alpha)=\emptyset$. Since $[t,t^{\prime}]\cap{\mathcal{S}}(\alpha)=\emptyset$, we have $D_{t}=D_{\tau}$ for all $\tau\in[t,t^{\prime}]$. Then, by the Smith inequality, (4.11), again, applied $k$ times to $I=[0,\tau]$, we have
@@
for all $k\in{\mathbb{N}}$. Here $p^{k}\tau$ ranges through the interval $I_{k}=[p^{k}t,p^{k}t^{\prime}]$ as $\tau$ ranges through $[t,t^{\prime}]$. Now let $C_{l}\to\infty$ be as in Proposition 4.6. Since the sequence $C_{l}$ has bounded gap, for all sufficiently large $k$, the intersection of the sequence $\{C_{l}\}$ and $I_{k}$ is non-empty: $\{C_{l}\}\cap[p^{k}t,p^{k}t^{\prime}]\neq\emptyset$. Therefore, there exists $\tau\in[t,t^{\prime}]$ and $k\in{\mathbb{N}}$ and $l\in{\mathbb{N}}$ such that $p^{k}\tau=C_{l}$. By Proposition 4.6, $D_{C_{l}}\leq 1$. Combining these facts, we arrive at
@@
Observe that by (B2) the difference $\big{|}\operatorname{\hat{\mu}}(y)-\operatorname{\hat{\mu}}(x)\big{|}$ is a priori bounded from above, by say, $2n$. Here we use the fact that $|\mu_{\pm}-\operatorname{\hat{\mu}}|<n-1$ by (3.2). Set $\rho={\mathcal{A}}(x)/\operatorname{\hat{\mu}}(x)$ and $\rho^{\prime}={\mathcal{A}}(y)/\operatorname{\hat{\mu}}(y)$. Then a direct calculation shows that
$b-a>|\rho^{\prime}/\rho-1|\cdot a-2n\rho.$
@@
By Theorem 3.3 and Corollary 3.5, there exists a positive real sequence $C_{l}\to\infty$ and $i_{0}$-sequences of positive even integers $d_{il}\to\infty$ as $l\to\infty$ such that for every iterated orbit $x_{ij}^{k}$ we have
@@
Indeed, in the setting of Theorem 3.3 and Corollary 3.5 we have $k\leq k_{ijl}$ in (4.13) and $k>k_{ijl}$ in (4.14). Here both action inequalities are strict, since the sequence $C_{l}$ can be chosen outside the action spectrum ${\mathcal{S}}(\alpha)$.
@@
Let $C_{l}\to\infty$ and $d_{il}\to\infty$ be as above. There exists a constant $K>0$ such that once $C_{l}>K$ we have
@@
In particular, $\operatorname{SH}^{C_{l}}(W)$ is supported in degrees of the same parity as $n$ since all $d_{il}$ are even.
@@
(Here we are using the fact that $i=i^{\prime}$, i.e., the orbits are in the same cluster.) Due to (B2), we have equality in both of these inequalities and the index difference is exactly $2$. Therefore, by (B3), $\deg(I)=\mu_{+}\big{(}x_{ij}^{k}\big{)}+1=d_{il}+n$. ∎
@@
At the same time, the left-hand side is equal $(-1)^{n}D$ by Lemma 4.9. Indeed, each term on the left is either 00 when $m\neq d_{il}+n$ or $(-1)^{n}\dim\operatorname{SH}^{C_{l}}_{m}(W)$ when $m=d_{il}+n$ since $d_{il}$ is even. Thus

### v3

@@ 4.1.1. Proof of Theorem A
Next, let $x_{0}:=x,\,x_{1},\ldots,x_{r}$ be all prime closed orbits which are eventually $\mathbb{Q}$}-visible, i.e., $x_{j}^{k}\in{\mathcal{P}}$ for some $k\in{\mathbb{N}}$. Note that the orbits $x_{j\geq 1}$ might be $\mathbb{Q}$}-invisible and/or have action below ${\mathcal{A}}(x)$. In fact, $x_{j}^{s}$ with $j\geq 1$ is necessarily $\mathbb{Q}$}-invisible when ${\mathcal{A}}\big(x_{j}^{s}\big)\leq{\mathcal{A}}(x)$. For each $j\geq 1$, we denote by $s_{j}$ the largest such $s$, setting $s_{j}=0$ if ${\mathcal{A}}(x_{j})>{\mathcal{A}}(x)$. Our goal is to show that there are exactly $n$ virtually $\mathbb{Q}$}-visible orbits among the orbits $x_{j}$.
@@
Indeed, then the only closed orbits with degree in $\mathcal{I}$} must be from $\Gamma_{0}$. By (4.1), since $N$ is divisible by all $N(x_{j})$, an orbit $x_{j}^{k_{j}}$ is $\mathbb{Q}$}-visible if and only if $x_{j}$ is virtually $\mathbb{Q}$}-visible. There are exactly $n$ integer spots in $\mathcal{I}$} of parity $n+1$, which must all be filled. Different $\mathbb{Q}$}-visible orbits have different degrees by Theorem E. Therefore, there are exactly $n$ virtually $\mathbb{Q}$}-visible prime orbits.
@@
Here in the first equality we used Lemma 2.8. The second equality follows from the fact that $\operatorname{CH}(x;{\mathbb{Q}})$ is concentrated in degree $n+1$ by Corollary D. As a consequence, $x^{k}\in{\mathcal{P}}$. We will need the following observation.
@@
We already know that $\operatorname{CH}(x;{\mathbb{Q}})$ and $\operatorname{CH}\big(x^{k};{\mathbb{Q}}\big)$ are supported in one degree and one-dimensional. Hence we only need to determine the degree shift $L$. Let $\varphi$ be the Poincaré return map of $x$. Then the return map of $x^{l}$ is $\varphi^{l}$ for all $l\in{\mathbb{N}}$. By (2.10) and (2.17),
$\dim\operatorname{CH}\big(x^{l};{\mathbb{Q}}\big)\leq\dim\operatorname{HF}\big(\varphi^{l};{\mathbb{Q}}\big),$
again for all $l\in{\mathbb{N}}$, and
$\dim\operatorname{CH}\big(x^{k};{\mathbb{Q}}\big)=1$
when $k$ is admissible. Moreover, $\operatorname{CH}(x;{\mathbb{Q}})=\operatorname{HF}(\varphi;{\mathbb{Q}})$ due to [Fe20, Thm. 1.2]; see also [Fe21]. Furthermore, as readily follows from [GG10, Thm. 1.1] (and its proof),
@@
since $k$ is admissible. Here the shift $L$ is exactly as in (D1) and (D2) depending on whether $x$ is totally degenerate or not. (Indeed, in the former case $\operatorname{HF}\big(\varphi^{k};{\mathbb{Q}}\big)=\operatorname{HF}(\psi\varphi;{\mathbb{Q}})$ for a loop $\psi$ of local Hamiltonian diffeomorphisms with $\operatorname{\hat{\mu}}(\psi)=(k-1)\operatorname{\hat{\mu}}(x)$ by [GG10, Claim 4.1]. Hence, $\operatorname{HF}\big(\varphi^{k};{\mathbb{Q}}\big)=\operatorname{HF}(\varphi;{\mathbb{Q}})[-(k-1)\operatorname{\hat{\mu}}(x)]$ by [GG10, (LF6)]. The latter case follows from [GG10, Sect. 4.5].) Combining these facts we conclude that $\operatorname{HF}\big(\varphi^{k};{\mathbb{Q}}\big)$ is one-dimensional and supported in $\deg(x^{k})$, and thus $\deg(x^{k})=\deg(x)+L$. ∎
@@
$\begin{split}\deg\big(x^{k_{0}-1}\big)&=(k_{0}-1)\operatorname{\hat{\mu}}(x)+q\\ &=\operatorname{\hat{\mu}}\big(x^{k_{0}}\big)-\operatorname{\hat{\mu}}(x)+q\\ &\leq\operatorname{\hat{\mu}}\big(x^{k_{0}}\big)-(n+1)\\ &\leq d+\eta-(n+1)\\ &<d-n.\end{split}$
In the last inequality, we use the fact that $\eta$ is small and, in particular, $\eta<1$. (In fact, in this case $d=\operatorname{\hat{\mu}}\big(x^{k_{0}}\big)$ because $x$ is totally degenerate, and hence $\operatorname{\hat{\mu}}\big(x^{k_{0}}\big)\in{\mathbb{Z}}$.) Thus $\deg\big(x^{k}\big)<d-n$ proving (4.5) when $x$ is totally degenerate.
@@
This guarantees, again by (4.3), that $x_{j}^{k_{j}-s}$ is an admissible odd iterate of $x_{j}^{s}$. Finally, by Lemma 2.8,
@@
This completes the proof of (4.4) and Theorem 4.1. ∎
@@ 4.1.2. Proof of Theorem B
Here is a detailed argument. Recall that when $M$ is centrally symmetric, there are two types of closed Reeb orbits in $M$: symmetric and asymmetric. The former are the closed orbits $x$ such that x(t)=x(t+T/2)-x(t)=x(t+T/2) where $T={\mathcal{A}}(x)$ is the period of $x$, and the latter are the orbits $x$ such that x-x and $x$ are geometrically distinct. Clearly, asymmetric orbits come in pairs, (x,x)(x,-x), with both $x$ and x-x having the same action and the same index. Thus, by Corollary D, the flow has no asymmetric orbits.
To finish the proof it suffices to show that all symmetric closed orbits are non-alternating. This is a standard observation going back to [LLZ]; below we closely follow [GiM].
Consider the homogeneous of degree two Hamiltonian $H$ on ${\mathbb{R}}^{2n}$ with $\{H=1\}=M$. The flow of $H$ on $M$ agrees with the Reeb flow on $M$. Let $\Psi^{t}$ be the linearized flow of $H$ in ${\mathbb{R}}^{2n}$ along a closed symmetric orbit $x$ in $M$ with period $T$. Using the constant trivialization $T{\mathbb{R}}^{2n}\cong{\mathbb{R}}^{2n}\times{\mathbb{R}}^{2n}$ we can view $\Psi^{t}$ as a path in $\operatorname{Sp}(2n)$. Then $\Psi^{T}=(\Psi^{T/2})^{2}$, since the trivialization and $H$ are both centrally symmetric, and therefore, $\Psi^{T}$ is non-alternating. To be more precise, $\Psi^{T}$ has an even number of eigenvalues in the range (1,0)(-1,0) and 1-1 is not an eigenvalue since the Reeb flow is non-degenerate. Fix a point $z$ on $x$ and consider the decomposition $T_{z}{\mathbb{R}}^{2n}=\xi_{z}\oplus{\mathbb{R}}^{2}$, where $\xi_{z}\subset T_{z}M$ is the contact plane and the second factor is the span of the Reeb vector field vector at $z$ and the radial direction. The decomposition is invariant under $\Psi^{T}\colon T_{z}{\mathbb{R}}^{2n}\to T_{z}{\mathbb{R}}^{2n}$, which then decomposes as $\Phi\oplus{\mathit{i}d}$, where $\Phi\colon\xi_{z}\to\xi_{z}$ is the linearized Reeb flow in time $T$. It follows that $\Phi$ and $\Psi^{T}$ have the same eigenvalues other than 1. Therefore, $\Phi$ and hence, by definition, $x$ are also non-alternating.
@@
It is illuminating to examine the role of symmetry, non-degeneracy and our other main results in the proof of Theorem B. The non-degeneracy condition is used to have non-alternating closed orbits defined and then, in some sense even more crucially, to guarantee that all iterates except the bad ones are $\mathbb{Q}$}-visible; cf. Examples 2.12 and 2.13. The lower bound $n$ on the number of non-alternating prime orbits follows in essence from [GK, LZ]; cf. Corollary 3.7. The genuinely new part of Theorem A is not needed here. However, we would have the same lower bound on the total number of prime closed orbits even without non-degeneracy by (M1). Likewise, if we replaced dynamical convexity by convexity or strong dynamical convexity in Theorem B, which are admittedly more restrictive conditions, by [LLZ] or [GiM]; cf. Corollary 3.8.
The symmetry condition together with the fact that distinct orbits have distinct actions (Corollary D) is central to the proof of Theorem B to rule out the existence of asymmetric closed orbits. This step is independent of the non-degeneracy requirement.
@@ 4.2. Proof of Theorem E
The proof relies on Theorem C and Corollary D, which are the central parts of the argument. Recall that $\mathcal{P}$} is the set of $\mathbb{Q}$}-visible closed orbits and $\mathcal{P}$$^{\prime}$ stands for $\mathcal{P}$} with the extra element $[W]$ of degree $n$ and action 0. We start with the proof of Part (O1), noting first that the existence of $x\in{\mathcal{P}}$ with $\deg(x)=m$ for any $m\in n-1+2{\mathbb{N}}$ is an immediate consequence of (2.7) and (2.22).
@@
It immediately follows from (4.10) and (4.11) that the sequence $\deg(I_{i})$ has constant parity. Since $\deg(I_{0})=n$, all $\deg(I_{i})$ have parity $n$.
@@
Furthermore, if we had $y\in{\mathcal{P}}$ with $q:=\deg(y)\in n+2{\mathbb{N}}$ we would have at least two orbits in $\mathcal{P}$} of degree q1q-1 or $q+1$ due to (2.22) or, more specifically, (2.23) and Theorem C. This is impossible since, as we have shown, for every $m\in n-1+2{\mathbb{N}}$ an orbit $x\in{\mathcal{P}}$ with $\deg(x)=m$ is unique. This completes the proof of the statement that the degree map $\deg$ on $\mathcal{P}$} takes values in $n-1+2{\mathbb{N}}$ and is a bijection.
@@
The statement that the ordering of the orbits in $\mathcal{P}$} (or the bars in $\mathcal{B}$}) by the action agrees with the ordering by the index is simply a reformulation of the fact that the sequence $\deg(x_{i})$ is strictly increasing. (Recall that the sequence $x_{i}$ is originally ordered by the action.) Alternatively, this statement is also a consequence of [GG20, Thm. 1.3]. This concludes the proof of Part (O1).
@@
When the flow is non-degenerate, Part (O1) has an analogue for any field $\mathbb{F}$} with $\deg(x)$ replaced by $\mu(x)$, although the statement is then less precise. Namely, the sequences $\mu(x_{i})$ and $\deg(I_{i})$ are still increasing although now not necessarily strictly. However, for every $m\in n-1+2{\mathbb{N}}$ we still have exactly one $\mathbb{F}$}-visible closed orbit $x$ with $\mu(x)=m$ and all $\deg(I_{i})$ have parity $n$. The key difference is that in this case we cannot rule out the existence of closed $\mathbb{F}$}-visible orbits $x$ with $\mu(x)$ of parity $n$ and intervals in the sequences $\deg(x_{i})$ and $\deg(I_{i})$ where these sequences are constant. The proof is a subset of the proof of Part (O1) combined with the fact that $\operatorname{supp}\operatorname{SH}(x;{\mathbb{F}})=\{\mu(x),\mu(x)+1\}$ for any field $\mathbb{F}$} when $x$ is non-degenerate, i.e., the second part of Corollary D holds automatically. We do not know to what degree Part (O2) extends to this case.
@@ 4.3.1. Generalities
This proposition is proved in Section 4.3.2. Here we only point out that this is the only part of the proof of Theorem C where the finiteness and dynamical convexity conditions are used. The rest of the argument is independent of these requirements. Note also that the sequence $C_{l}$ can be taken independent of the field $\mathbb{F}$}. However, we do not need this fact.
@@
Thus, fix $t\in{\mathbb{R}}_{+}\setminus{\mathcal{S}}(\alpha)$ and let $t^{\prime}>t$ be such that $t,t^{\prime}]\cap{\mathcal{S}}(\alpha)=\emptyset$. Since $t,t^{\prime}]\cap{\mathcal{S}}(\alpha)=\emptyset$, we have $D_{t}=D_{\tau}$ for all $\tau\in[t,t^{\prime}]$. Then, by the Smith inequality, (4.12), again, applied $k$ times to $I=[0,\tau]$, we have
@@
for all $k\in{\mathbb{N}}$. Here $p^{k}\tau$ ranges through the interval $I_{k}=[p^{k}t,p^{k}t^{\prime}]$ as $\tau$ ranges through $t,t^{\prime}]$. Now let $C_{l}\to\infty$ be as in Proposition 4.8. Since the sequence $C_{l}$ has bounded gap, for all sufficiently large $k$, the intersection of the sequence $\{C_{l}\}$ and $I_{k}$ is non-empty: $\{C_{l}\}\cap[p^{k}t,p^{k}t^{\prime}]\neq\emptyset$. Therefore, there exists $\tau\in[t,t^{\prime}]$ and $k\in{\mathbb{N}}$ and $l\in{\mathbb{N}}$ such that $p^{k}\tau=C_{l}$. By Proposition 4.8, $D_{C_{l}}\leq 1$. Combining these facts, we arrive at
@@
Observe that by (B2) the difference $\big|\operatorname{\hat{\mu}}(y)-\operatorname{\hat{\mu}}(x)\big|$ is a priori bounded from above by $2n$. Here we use the fact that $|\mu_{\pm}-\operatorname{\hat{\mu}}|\leq n-1$ by (3.2). Set $\rho={\mathcal{A}}(x)/\operatorname{\hat{\mu}}(x)$ and $\rho^{\prime}={\mathcal{A}}(y)/\operatorname{\hat{\mu}}(y)$ and recall that $b={\mathcal{A}}(y)$ and $a={\mathcal{A}}(x)$. Then a direct calculation shows that
$b-a\geq|\rho^{\prime}/\rho-1|\cdot b-2n\rho.$
@@
By Theorem 3.3 and Corollary 3.5, there exists a positive real sequence $C_{l}\to\infty$ and for each $i=1,\ldots,i_{0}$ there is a sequence of positive even integers $d_{il}\to\infty$ as $l\to\infty$ such that for every iterated orbit $x_{ij}^{k}$ we have
@@
Indeed, in the setting of Theorem 3.3 and Corollary 3.5 for a recurrence event $k_{ijl}$ we have $k\leq k_{ijl}$ in (4.14) and $k>k_{ijl}$ in (4.15). Here both action inequalities are strict, since the sequence $C_{l}$ can be chosen outside the action spectrum ${\mathcal{S}}(\alpha)$.
@@
Let $C_{l}\to\infty$ and $d_{il}\to\infty$ be as above, i.e., these sequences satisfy (4.14) and (4.15) and all $d_{il}$ are even. Then there exists a constant $K>0$ such that once $C_{l}>K$ we have
@@
In particular, $\operatorname{SH}^{C_{l}}(W)$ is supported in degrees of the same parity as $n$.
@@
(Here we are using the fact that $i=i^{\prime}$, i.e., the orbits are in the same cluster.) Due to (B2), we have equality in both of these inequalities and the index difference is exactly 22. Therefore, by (B3), $\deg(I)=\mu_{+}\big(x_{ij}^{k}\big)+1=d_{il}+n$. ∎
@@
At the same time, the left-hand side is equal to $(-1)^{n}D$ by Lemma 4.11. Indeed, each term on the left is either 0 when $m\neq d_{il}+n$ or $(-1)^{n}\dim\operatorname{SH}^{C_{l}}_{m}(W)$ when $m=d_{il}+n$ since $d_{il}$ is even. Thus